An athlete is running in five races and in each race he has a 70% chance of winning. What is the probability that he will win at least two races?


1
Expert's answer
2021-06-11T10:10:58-0400

Let "X=" the number of winned races: "X\sim Bin(n, p)"

Given "n=5, p=0.7"


"P(X\geq2)=1-P(X=0)-P(X=1)"

"=1-\dbinom{5}{0}(0.7)^0(1-0.7)^{5-0}-\dbinom{5}{1}(0.7)^1(1-0.7)^{5-1}"

"=1-(0.3)^5-5(0.7)(0.3)^4"

"=1-3.8(0.3)^4=0.96922"

The probability that he will win at least two races is 0.96922.